LINE PERPENDICULAR TO -4X+Y=-25 AND CONTAINS (4,-10)
Answer by Cromlix(4381)

Sort -4x + y = -25
into y = mx + c form.
y = 4x - 25
Lines that are perpendicular to
one another have slopes (m)
that multiply together to give -1
m1 * m2 = -1
This line has an 'm' (slope) = 4
4 * m2 = -1
m2 = -1/4
Using coordinates (4, -10)
in equation y - b = m(x - a)
a = 4, b = -10 and m = -1/4
y - (-10) = -1/4(x - 4)
y + 10 = -1/4x + 4/4
y = -1/4x + 1 - 10
y = -1/4x - 9
OR
multiply thro' by 4
4y = -x - 36
Hope this helps.
:-)